# Drinfeld rational fractions for affine Kac–Moody quantum symmetric pairs

### v5

@@ 4.4. Proof of the factorization theorem
(4.49) $(1-q^{-2}Cz^{2})\boldsymbol{\Theta}(z)\equiv$ $\ \frac{(K-CK^{-1}z^{2})\boldsymbol{\phi}(z^{-1})}{q-q^{-1}}$
(4.50) $\ -q^{2}c_{0}z\boldsymbol{\phi}(z^{-1})[\mathbb{A},(1+q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)x_{1}^{-}]\quad\mod{U_{+}}[\negthinspace[z]\negthinspace].$
@@
(4.63) $\equiv$ $\ \frac{(K-CK^{-1}z^{2})\boldsymbol{\phi}(z^{-1})+(Cz^{2}-1)}{c_{1}^{-1}Cz^{2}(q-q^{-1})}.$
@@
(4.64) $(z^{-1}A_{-1}^{(-1)}-q^{-2}A_{0}^{(-1)})\boldsymbol{A}^{(1)}_{+}(z)=$ $\ (z^{-1}A_{-1}^{(-1)}-q^{-2}A_{0}^{(-1)})\boldsymbol{\phi}(z^{-1})\mathbb{A}$
(4.65) $=$ $\ (x_{1}^{-}z^{-1}-q^{-2}x_{0}^{-})\boldsymbol{\phi}(z^{-1})\mathbb{A}$
(4.66) $=$ $\ \boldsymbol{\phi}(z^{-1})(q^{-2}x_{1}^{-}z^{-1}-x_{0}^{-})\mathbb{A},$
@@
$\boldsymbol{A}^{(1)}_{+}(z)(-q^{-2}z^{-1}A_{-1}^{(-1)}+A_{0}^{(-1)})=\boldsymbol{\phi}(z^{-1})\mathbb{A}(-q^{-2}x^{-}_{1}z^{-1}+x^{-}_{0}).$
@@
(4.68) $=$ $\ -q^{-2}z^{-1}\boldsymbol{\phi}(z^{-1})[\mathbb{A},(1+q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)x_{1}^{-}].$
@@
(4.79) $\boldsymbol{\Theta}(z)\equiv\frac{1-Cz^{2}}{(q-q^{-1})(1-q^{-2}Cz^{2})}\boldsymbol{\phi}(z^{-1})\boldsymbol{\psi}(Cz)\mod{U_{+}}[\negthinspace[z]\negthinspace].$
@@
(4.80) $-q^{2}c_{0}z\boldsymbol{\phi}(z^{-1})[\mathbb{A},(1\ +$ $\ q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)x_{1}^{-}]=-q^{2}c_{0}z\boldsymbol{\phi}(z^{-1})[(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)\mathbb{A},x_{1}^{-}]$
(4.81) $=$ $\ Cz\boldsymbol{\phi}(z^{-1})[(1-\operatorname{ad}_{\overline{h}_{-1}}z)\boldsymbol{x}^{+}_{\geq 0}(Cz),x^{-}_{1}]$
(4.82) $=$ $\ Cz\boldsymbol{\phi}(z^{-1})[\boldsymbol{x}^{+}_{\geq 0}(Cz),(1+\operatorname{ad}_{\overline{h}_{-1}}z)x^{-}_{1}]$
(4.83) $=$ $\ \frac{Cz\boldsymbol{\phi}(z^{-1})}{q-q^{-1}}\big{(}(Cz)^{-1}\boldsymbol{\psi}_{\geq 1}(Cz)-z\boldsymbol{\psi}(Cz)+zK^{-1}\big{)}$
(4.84) $=$ $\ \frac{\boldsymbol{\phi}(z^{-1})}{q-q^{-1}}\big{(}(1-Cz^{2})\boldsymbol{\psi}(Cz)-K+Cz^{2}K^{-1}\big{)}.$
@@
(4.86) $\boldsymbol{\grave{\Theta}}(z)\equiv\boldsymbol{\phi}(z^{-1})\boldsymbol{\psi}(Cz){\mod U_{+}}[\negthinspace[z]\negthinspace].$
@@ 5.1. Spectra of restricted representations
Let $\mathcal{H}$ (resp. $\mathcal{H}_{\mathcal{O}_{q}}$) be the commutative subalgebra of $U_{q}L\mathfrak{sl}_{2}$ (resp. $\mathcal{O}_{q}$) generated by elements $\phi_{i},\psi_{i}$ (resp. $\Theta_{i}$). Recall that every irreducible finite dimensional $U_{q}L\mathfrak{sl}_{2}$-module $V$ can be endowed with a ${\mathbb{Z}}_{\leq 0}$-grading (compatible with the ${\mathbb{Z}}$-grading on $U_{q}L\mathfrak{sl}_{2}$) such that the highest weight vector is in degree zero. Indeed, this is clear, in the case of evaluation representations, from their explicit description, and follows in the general case from the fact that every finite dimensional irreducible module is obtained as a tensor product of evaluation representations [CP91, Theorem 4.11]. This grading also induces a (vector space) filtration with $F_{i}(V)=\{v\in V\mid\deg v\geq i\}$. Since this filtration is preserved by the action of $\mathcal{H}$, we can consider the associated graded $\operatorname{gr}(V)$ as a $\mathcal{H}$-module.
The $\mathcal{O}_{q}$-module $\eta_{\mathbf{c},0}^{*}(V)$ has a decomposition into generalized $\mathcal{H}_{\mathcal{O}_{q}}$-eigenspaces. Let $w$ be generalized $\mathcal{H}_{\mathcal{O}_{q}}$-eigenvector and $\overline{w}$ its image in $\operatorname{gr}(V)$. Corollary 4.11 implies that $\overline{w}$ is an $\mathcal{H}$-eigenvector, and that the $\mathcal{H}_{\mathcal{O}_{q}}$-eigenvalues of $w$ can be computed using the $\mathcal{H}$-eigenvalues of $\overline{w}$. More precisely, in terms of generating series, the eigenvalue of $\boldsymbol{\grave{\Theta}}(z)$ on $w$ is the same as the $\boldsymbol{\phi}(z^{-1})\boldsymbol{\psi}(Cz)$-eigenvalue on $\overline{w}$. This fact has many interesting consequences.
@@ 5.2. Frenkel–Reshetikhin-type theorem
(5.1) $\boldsymbol{\psi}(z)\cdot\overline{w}=\boldsymbol{d}_{w}^{+}(z)\overline{w},\quad\boldsymbol{\phi}(z)\cdot\overline{w}=\boldsymbol{d}_{w}^{-}(z)\overline{w},\quad\boldsymbol{\grave{\Theta}}(z)\cdot w=\boldsymbol{D}_{w}(z)w.$
@@ 5.3. Rational fractions
(5.12) $=$ $\ (qC^{-\frac{1}{2}}z^{-1})^{\deg\mathbf{Q}}\frac{\mathbf{Q}(q^{-1}z)}{\mathbf{Q}^{\dagger}(qC^{-1}z^{-1})},$
@@
(5.15) $=$ $\ (qC^{-\frac{1}{2}}z^{-1})^{\deg\mathbf{Q}}\frac{\mathbf{Q}(q^{-1}z)}{\mathbf{Q}^{\dagger}(qC^{-1}z^{-1})},$
@@ 7.1. Coproduct formula for $\boldsymbol{A}_{+}(z)$
(7.8) $[Q_{4},S^{r}_{1}]+[Q_{2},S^{r}_{2}]=$ $\ -CA_{-1}\otimes(q-q^{-1})[x^{+}_{0},x^{-}_{-r}]+[\overline{H}_{1},A_{0}]\otimes\phi_{-r}$
(7.9) $=$ $\ ([\overline{H}_{1},A_{0}]+CA_{-1})\otimes\phi_{-r}=A_{1}\otimes\phi_{-r}.$
@@
where $\widetilde{\xi}_{\mathbf{c},\mathbf{s}}$ is as in (2.23). From now on all the elements on the LHS of the tensor symbol are understood as images under $\xi_{\mathbf{c},\mathbf{s}}$. For example, $A_{0}\otimes\phi_{-r}=\xi_{\mathbf{c},\mathbf{s}}(A_{0})\otimes\phi_{-r}$, etc. In particular, any two elements on the LHS of the tensor symbol commute. Abbreviate
@@
(7.10) $S^{r}_{4}=\sum_{\begin{subarray}{c}i+j+k=r;\\ i\geq 1;j,k\geq 0\end{subarray}}A_{k}\otimes\kappa\mathbbm{h}_{i-1}[x^{+}_{-1},\phi_{-j}],$
@@
(7.13) $\overline{\Delta}_{\mathbf{c},\mathbf{s}}(\boldsymbol{A}_{+}(z))\equiv$ $\ 1\otimes\boldsymbol{A}_{+}(z)+\boldsymbol{A}_{+}(z)\otimes\boldsymbol{\phi}(z^{-1})+\boldsymbol{A}_{+}(z)\otimes\kappa\Gamma,$
@@
$\Gamma=(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)^{-1}(1-C\operatorname{ad}_{\overline{h}_{1}}z)^{-1}(q^{2}-1)\boldsymbol{\phi}(z^{-1})x^{+}_{-1}z.$
@@
$\sum_{r\geq 0}S^{r}_{1}=1\otimes\boldsymbol{A}_{+}(z),\qquad\sum_{r\geq 0}S^{r}_{2+3}=\boldsymbol{A}_{+}(z)\otimes\boldsymbol{\phi}(z^{-1}).$
@@
(7.14) $\sum_{r\geq 1}S^{r}_{4}z^{r}=$ $\ \boldsymbol{A}_{+}(z)\otimes z\kappa\Omega^{-1}[x^{+}_{-1},\boldsymbol{\phi}(z^{-1})]=\boldsymbol{A}_{+}(z)\otimes\kappa\Gamma,$
@@
so that $\Gamma=\boldsymbol{\phi}(z^{-1})\widetilde{\Gamma}$.
@@ 7.2. Coproduct formula for $\boldsymbol{\grave{\Theta}}(z)$
(7.23) $\equiv$ $\ z^{2}c_{1}^{-1}C\boldsymbol{A}_{+}(z)\otimes\left(z^{-1}[x^{-}_{1},\boldsymbol{\phi}(z^{-1})]_{q^{-2}}-q^{-2}[x^{-}_{0},\boldsymbol{\phi}(z^{-1})]_{q^{2}}\right)$
@@
(7.26) $\equiv$ $\ z^{2}c_{1}^{-1}C(1-q^{-2})\boldsymbol{A}_{+}(z)\otimes\left(z^{-1}\xi_{\mathbf{c},\mathbf{s}}(A_{-1})K+\xi_{\mathbf{c},\mathbf{s}}(A_{0})K^{-1}\right)\boldsymbol{\phi}(z^{-1}).$
@@
(7.28) $\equiv$ $\ z^{2}c_{1}^{-1}C(1-q^{-2})\boldsymbol{A}_{+}(z)\otimes\left(z^{-1}\xi_{\mathbf{c},\mathbf{s}}(A_{-1})K+\xi_{\mathbf{c},\mathbf{s}}(A_{0})K^{-1}\right)\boldsymbol{\phi}(z^{-1}).$
@@
(7.31) $\equiv$ $\ -q^{-2}zc_{1}^{-1}C\boldsymbol{A}_{+}(z)\otimes\boldsymbol{\phi}(z^{-1})[\kappa\widetilde{\Gamma},(1+q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)x^{-}_{1}]$
(7.32) $\equiv$ $\ -q^{-2}zc_{1}^{-1}C\boldsymbol{A}_{+}(z)\otimes\boldsymbol{\phi}(z^{-1})[(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)\kappa\widetilde{\Gamma},x^{-}_{1}]$
(7.33) $\equiv$ $\ q^{-1}z^{2}(q-q^{-1})^{2}c_{1}^{-1}C\boldsymbol{A}_{+}(z)\otimes\boldsymbol{\phi}(z^{-1})\cdot$
@@
(7.38) $=$ $\ z^{-1}(\xi_{\mathbf{c},\mathbf{s}}(A_{-1})+\xi_{\mathbf{c},\mathbf{s}}(A_{0})z)\boldsymbol{\psi}_{\geq 1}(Cz)+\xi_{\mathbf{c},\mathbf{s}}(A_{0})(K-K^{-1})$
(7.39) $=$ $\ (\xi_{\mathbf{c},\mathbf{s}}(A_{-1})z^{-1}+\xi_{\mathbf{c},\mathbf{s}}(A_{0}))\boldsymbol{\psi}(Cz)-\xi_{\mathbf{c},\mathbf{s}}(A_{-1})z^{-1}K-\xi_{\mathbf{c},\mathbf{s}}(A_{0})K^{-1}.$
@@
(7.41) $\equiv$ $\ q^{-1}z^{2}(q-q^{-1})c_{1}^{-1}C\boldsymbol{A}_{+}(z)\otimes\boldsymbol{\phi}(z^{-1})\cdot$
(7.42) $\{(\xi_{\mathbf{c},\mathbf{s}}(A_{-1})z^{-1}+\xi_{\mathbf{c},\mathbf{s}}(A_{0}))\boldsymbol{\psi}(Cz)$
@@
(7.45) $+z(1-q^{-2})c_{1}^{-1}C(A_{-1}+A_{0}z)\boldsymbol{A}_{+}(z)\otimes\boldsymbol{\phi}(z^{-1})\boldsymbol{\psi}(Cz).$

### v6

@@ 4.4. Proof of the factorization theorem
(4.49) $(1-q^{-2}Cz^{2})\boldsymbol{\Theta}(z)\equiv$ $\ \frac{(K-CK^{-1}z^{2})\boldsymbol{\phi}^{-}(z^{-1})}{q-q^{-1}}$
(4.50) $\ -q^{2}c_{0}z\boldsymbol{\phi}^{-}(z^{-1})[\mathbb{A},(1+q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)x_{1}^{-}]\quad\mod{U_{+}}[\negthinspace[z]\negthinspace].$
@@
(4.63) $\equiv$ $\ \frac{(K-CK^{-1}z^{2})\boldsymbol{\phi}^{-}(z^{-1})+(Cz^{2}-1)}{c_{1}^{-1}Cz^{2}(q-q^{-1})}.$
@@
(4.64) $(z^{-1}A_{-1}^{(-1)}-q^{-2}A_{0}^{(-1)})\boldsymbol{A}^{(1)}_{+}(z)=$ $\ (z^{-1}A_{-1}^{(-1)}-q^{-2}A_{0}^{(-1)})\boldsymbol{\phi}^{-}(z^{-1})\mathbb{A}$
(4.65) $=$ $\ (x_{1}^{-}z^{-1}-q^{-2}x_{0}^{-})\boldsymbol{\phi}^{-}(z^{-1})\mathbb{A}$
(4.66) $=$ $\ \boldsymbol{\phi}^{-}(z^{-1})(q^{-2}x_{1}^{-}z^{-1}-x_{0}^{-})\mathbb{A},$
@@
$\boldsymbol{A}^{(1)}_{+}(z)(-q^{-2}z^{-1}A_{-1}^{(-1)}+A_{0}^{(-1)})=\boldsymbol{\phi}^{-}(z^{-1})\mathbb{A}(-q^{-2}x^{-}_{1}z^{-1}+x^{-}_{0}).$
@@
(4.68) $=$ $\ -q^{-2}z^{-1}\boldsymbol{\phi}^{-}(z^{-1})[\mathbb{A},(1+q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)x_{1}^{-}].$
@@
(4.79) $\boldsymbol{\Theta}(z)\equiv\frac{1-Cz^{2}}{(q-q^{-1})(1-q^{-2}Cz^{2})}\boldsymbol{\phi}^{-}(z^{-1})\boldsymbol{\phi}^{+}(Cz)\mod{U_{+}}[\negthinspace[z]\negthinspace].$
@@
(4.80) $-q^{2}c_{0}z\boldsymbol{\phi}^{-}(z^{-1})[\mathbb{A},(1\ +$ $\ q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)x_{1}^{-}]=-q^{2}c_{0}z\boldsymbol{\phi}^{-}(z^{-1})[(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)\mathbb{A},x_{1}^{-}]$
(4.81) $=$ $\ Cz\boldsymbol{\phi}^{-}(z^{-1})[(1-\operatorname{ad}_{\overline{h}_{-1}}z)\boldsymbol{x}^{+}_{\geq 0}(Cz),x^{-}_{1}]$
(4.82) $=$ $\ Cz\boldsymbol{\phi}^{-}(z^{-1})[\boldsymbol{x}^{+}_{\geq 0}(Cz),(1+\operatorname{ad}_{\overline{h}_{-1}}z)x^{-}_{1}]$
(4.83) $=$ $\ \frac{Cz\boldsymbol{\phi}^{-}(z^{-1})}{q-q^{-1}}\big{(}(Cz)^{-1}\boldsymbol{\phi}^{+}_{\geq 1}(Cz)-z\boldsymbol{\phi}^{+}(Cz)+zK^{-1}\big{)}$
(4.84) $=$ $\ \frac{\boldsymbol{\phi}^{-}(z^{-1})}{q-q^{-1}}\big{(}(1-Cz^{2})\boldsymbol{\phi}^{+}(Cz)-K+Cz^{2}K^{-1}\big{)}.$
@@
(4.86) $\boldsymbol{\grave{\Theta}}(z)\equiv\boldsymbol{\phi}^{-}(z^{-1})\boldsymbol{\phi}^{+}(Cz){\mod U_{+}}[\negthinspace[z]\negthinspace].$
@@ 5.1. Spectra of restricted representations
Let $\mathcal{H}$ (resp. $\mathcal{H}_{\mathcal{O}_{q}}$) be the commutative subalgebra of $U_{q}L\mathfrak{sl}_{2}$ (resp. $\mathcal{O}_{q}$) generated by elements $\phi^{\pm}_{i}$ (resp. $\Theta_{i}$). Recall that every irreducible finite dimensional $U_{q}L\mathfrak{sl}_{2}$-module $V$ can be endowed with a ${\mathbb{Z}}_{\leq 0}$-grading (compatible with the ${\mathbb{Z}}$-grading on $U_{q}L\mathfrak{sl}_{2}$) such that the highest weight vector is in degree zero. Indeed, this is clear, in the case of evaluation representations, from their explicit description, and follows in the general case from the fact that every finite dimensional irreducible module is obtained as a tensor product of evaluation representations [CP91, Theorem 4.11]. This grading also induces a (vector space) filtration with $F_{i}(V)=\{v\in V\mid\deg v\geq i\}$. Since this filtration is preserved by the action of $\mathcal{H}$, we can consider the associated graded $\operatorname{gr}(V)$ as a $\mathcal{H}$-module.
The $\mathcal{O}_{q}$-module $\eta_{\mathbf{c},0}^{*}(V)$ has a decomposition into generalized $\mathcal{H}_{\mathcal{O}_{q}}$-eigenspaces. Let $w$ be generalized $\mathcal{H}_{\mathcal{O}_{q}}$-eigenvector and $\overline{w}$ its image in $\operatorname{gr}(V)$. Corollary 4.11 implies that $\overline{w}$ is a generalized $\mathcal{H}$-eigenvector, and that the $\mathcal{H}_{\mathcal{O}_{q}}$-eigenvalues of $w$ can be computed using the $\mathcal{H}$-eigenvalues of $\overline{w}$. More precisely, in terms of generating series, the eigenvalue of $\boldsymbol{\grave{\Theta}}(z)$ on $w$ is the same as the $\boldsymbol{\phi}^{-}(z^{-1})\boldsymbol{\phi}^{+}(Cz)$-eigenvalue on $\overline{w}$. This fact has many interesting consequences.
@@ 5.2. Frenkel–Reshetikhin-type theorem
(5.1) $\boldsymbol{\phi}^{+}(z)\cdot\overline{w}=\boldsymbol{d}_{w}^{+}(z)\overline{w},\qquad\boldsymbol{\phi}^{-}(z)\cdot\overline{w}=\boldsymbol{d}_{w}^{-}(z)\overline{w},\qquad\boldsymbol{\grave{\Theta}}(z)\cdot w=\boldsymbol{D}_{w}(z)w.$
@@ 5.3. Rational fractions
(5.12) $=$ $\ (qC^{-\frac{1}{2}}z^{-1})^{\deg\mathbf{Q}}\frac{\mathbf{Q}^{\dagger}(q^{-1}z)}{\mathbf{Q}^{\dagger}(qC^{-1}z^{-1})},$
@@
(5.15) $=$ $\ (qC^{-\frac{1}{2}}z^{-1})^{\deg\mathbf{Q}}\frac{\mathbf{Q}^{\dagger}(q^{-1}z)}{\mathbf{Q}^{\dagger}(qC^{-1}z^{-1})},$
@@ 7.1. Coproduct formula for $\boldsymbol{A}_{+}(z)$
(7.8) $[Q_{4},S^{r}_{1}]+[Q_{2},S^{r}_{2}]=$ $\ -CA_{-1}\otimes(q-q^{-1})[x^{+}_{0},x^{-}_{-r}]+[\overline{H}_{1},A_{0}]\otimes\phi^{-}_{-r}$
(7.9) $=$ $\ ([\overline{H}_{1},A_{0}]+CA_{-1})\otimes\phi^{-}_{-r}=A_{1}\otimes\phi^{-}_{-r}.$
@@
where $\widetilde{\xi}_{\mathbf{c},\mathbf{s}}$ is as in (2.24). From now on all the elements on the LHS of the tensor symbol are understood as images under $\xi_{\mathbf{c},\mathbf{s}}$. For example, $A_{0}\otimes\phi^{-}_{-r}=\xi_{\mathbf{c},\mathbf{s}}(A_{0})\otimes\phi^{-}_{-r}$, etc. In particular, any two elements on the LHS of the tensor symbol commute. Abbreviate
@@
(7.10) $S^{r}_{4}=\sum_{\begin{subarray}{c}i+j+k=r;\\ i\geq 1;j,k\geq 0\end{subarray}}A_{k}\otimes\kappa\mathbbm{h}_{i-1}[x^{+}_{-1},\phi_{-j}^{-}],$
@@
(7.13) $\overline{\Delta}_{\mathbf{c},\mathbf{s}}(\boldsymbol{A}_{+}(z))\equiv$ $\ 1\otimes\boldsymbol{A}_{+}(z)+\boldsymbol{A}_{+}(z)\otimes\boldsymbol{\phi}^{-}(z^{-1})+\boldsymbol{A}_{+}(z)\otimes\kappa\Gamma,$
@@
$\Gamma=(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)^{-1}(1-C\operatorname{ad}_{\overline{h}_{1}}z)^{-1}(q^{2}-1)\boldsymbol{\phi}^{-}(z^{-1})x^{+}_{-1}z.$
@@
$\sum_{r\geq 0}S^{r}_{1}=1\otimes\boldsymbol{A}_{+}(z),\qquad\sum_{r\geq 0}S^{r}_{2+3}=\boldsymbol{A}_{+}(z)\otimes\boldsymbol{\phi}^{-}(z^{-1}).$
@@
(7.14) $\sum_{r\geq 1}S^{r}_{4}z^{r}=$ $\ \boldsymbol{A}_{+}(z)\otimes z\kappa\Omega^{-1}[x^{+}_{-1},\boldsymbol{\phi}^{-}(z^{-1})]=\boldsymbol{A}_{+}(z)\otimes\kappa\Gamma,$
@@
so that $\Gamma=\boldsymbol{\phi}^{-}(z^{-1})\widetilde{\Gamma}$.
@@ 7.2. Coproduct formula for $\boldsymbol{\grave{\Theta}}(z)$
(7.23) $\equiv$ $\ z^{2}c_{1}^{-1}C\boldsymbol{A}_{+}(z)\otimes\left(z^{-1}[x^{-}_{1},\boldsymbol{\phi}^{-}(z^{-1})]_{q^{-2}}-q^{-2}[x^{-}_{0},\boldsymbol{\phi}^{-}(z^{-1})]_{q^{2}}\right)$
@@
(7.26) $\equiv$ $\ z^{2}c_{1}^{-1}C(1-q^{-2})\boldsymbol{A}_{+}(z)\otimes\left(z^{-1}\xi_{\mathbf{c},\mathbf{s}}(A_{-1})K+\xi_{\mathbf{c},\mathbf{s}}(A_{0})K^{-1}\right)\boldsymbol{\phi}^{-}(z^{-1}).$
@@
(7.28) $\equiv$ $\ z^{2}c_{1}^{-1}C(1-q^{-2})\boldsymbol{A}_{+}(z)\otimes\left(z^{-1}\xi_{\mathbf{c},\mathbf{s}}(A_{-1})K+\xi_{\mathbf{c},\mathbf{s}}(A_{0})K^{-1}\right)\boldsymbol{\phi}^{-}(z^{-1}).$
@@
(7.31) $\equiv$ $\ -q^{-2}zc_{1}^{-1}C\boldsymbol{A}_{+}(z)\otimes\boldsymbol{\phi}^{-}(z^{-1})[\kappa\widetilde{\Gamma},(1+q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)x^{-}_{1}]$
(7.32) $\equiv$ $\ -q^{-2}zc_{1}^{-1}C\boldsymbol{A}_{+}(z)\otimes\boldsymbol{\phi}^{-}(z^{-1})[(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)\kappa\widetilde{\Gamma},x^{-}_{1}]$
(7.33) $\equiv$ $\ q^{-1}z^{2}(q-q^{-1})^{2}c_{1}^{-1}C\boldsymbol{A}_{+}(z)\otimes\boldsymbol{\phi}^{-}(z^{-1})\cdot$
@@
(7.38) $=$ $\ z^{-1}(\xi_{\mathbf{c},\mathbf{s}}(A_{-1})+\xi_{\mathbf{c},\mathbf{s}}(A_{0})z)\boldsymbol{\phi}^{+}_{\geq 1}(Cz)+\xi_{\mathbf{c},\mathbf{s}}(A_{0})(K-K^{-1})$
(7.39) $=$ $\ (\xi_{\mathbf{c},\mathbf{s}}(A_{-1})z^{-1}+\xi_{\mathbf{c},\mathbf{s}}(A_{0}))\boldsymbol{\phi}^{+}(Cz)-\xi_{\mathbf{c},\mathbf{s}}(A_{-1})z^{-1}K-\xi_{\mathbf{c},\mathbf{s}}(A_{0})K^{-1}.$
@@
(7.41) $\equiv$ $\ q^{-1}z^{2}(q-q^{-1})c_{1}^{-1}C\boldsymbol{A}_{+}(z)\otimes\boldsymbol{\phi}^{-}(z^{-1})\cdot$
(7.42) $\{(\xi_{\mathbf{c},\mathbf{s}}(A_{-1})z^{-1}+\xi_{\mathbf{c},\mathbf{s}}(A_{0}))\boldsymbol{\phi}^{+}(Cz)$
@@
(7.45) $+z(1-q^{-2})c_{1}^{-1}C(A_{-1}+A_{0}z)\boldsymbol{A}_{+}(z)\otimes\boldsymbol{\phi}^{-}(z^{-1})\boldsymbol{\phi}^{+}(Cz).$